# The contact process on dynamical random trees with degree dependence

### v1

@@ 5.2 Properties of $(\mathbf{C},\mathbf{B})$ on $\mathcal{T}$
Let us now see the proof for $\lambda_{2}(\mathcal{T})$. Denote by
$\beta_{\mathcal{T}}(\lambda,v,\{\rho\}):=\liminf_{t\to\infty}\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(\rho\in\mathbf{C}_{t}).$
Appealing to the same arguments as in the first part, it is enough to prove that if  $\beta_{\mathcal{T}_{i}}(\lambda,v,\{x_{i}\})>0$ for some $i\in\{1,\dots,k\}$, then $\beta_{\mathcal{T}}(\lambda,v,\{\rho\})>0$, since we would get the inequality
$\mathbb{P}\big{(}\beta_{\mathcal{T}}(\lambda,v,\{\rho\})=0\big{)}\leq g\Big{(}\mathbb{P}\big{(}\beta_{\mathcal{T}}(\lambda,v,\{\rho\})=0\big{)}\Big{)},$
by an analogous procedure and therefore the claim follows. Let $i\in\{1,\dots,k\}$. Note that for $r$ such that $t-2r>0$, we have
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(\rho\in\mathbf{C}_{t})$$\geq\mathbb{P}^{\{\rho\}}(\rho\in\mathbf{C}_{t},\ x_{i}\in\mathbf{C}_{t-r},\ x_{i}\in\mathbf{C}_{r})$
$=\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(\rho\in\mathbf{C}_{t}\ |\ x_{i}\in\mathbf{C}_{r},\ x_{i}\in\mathbf{C}_{t-r})\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(x_{i}\in\mathbf{C}_{t-r}\ |\ x_{i}\in\mathbf{C}_{r})\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(x_{i}\in\mathbf{C}_{r})$
$\geq\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(\rho\in\mathbf{C}_{t}\ |\ x_{i}\in\mathbf{C}_{t-r})\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(x_{i}\in\mathbf{C}_{t-r}\ |\ x_{i}\in\mathbf{C}_{r})\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(x_{i}\in\mathbf{C}_{r}),$
where in the last inequality we have used monotonicity.
Now using again monotonicity of the contact processes, we get
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(\rho\in\mathbf{C}_{t}\ |\ x_{i}\in\mathbf{C}_{t-r})\geq\mathbb{P}_{\mathcal{T}}^{\{x_{i}\}}(\rho\in\mathbf{C}_{r}).$
and
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(x_{i}\in\mathbf{C}_{t-r}\ |\ x_{i}\in\mathbf{C}_{r})\geq\mathbb{P}_{\mathcal{T}_{i}}(x_{i}\in\mathbf{C}_{t-r}\ |\ x_{i}\in\mathbf{C}_{r})\geq\mathbb{P}_{\mathcal{T}_{i}}^{\{x_{i}\}}(x_{i}\in\mathbf{C}_{t-2r}),$
where the first inequality holds since switching from  $\mathcal{T}$ to  $\mathcal{T}_{i}$ makes the probability smaller since we are considering fewer neighbors. It follows that
$\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(\rho\in\mathbf{C}_{t})$ $\geq\mathbb{P}_{\mathcal{T}}^{\{x_{i}\}}(\rho\in\mathbf{C}_{r})\mathbb{P}_{\mathcal{T}_{i}}^{\{x_{i}\}}(x_{i}\in\mathbf{C}_{t-2r})\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(x_{i}\in\mathbf{C}_{r}).$ (5.3)
In addition, we observe that
$\mathbb{P}_{\mathcal{T}}^{\{x_{i}\}}(\rho\in\mathbf{C}_{r})\geq p(D_{\rho},D_{x_{i}})e^{-v(D_{\rho},D_{x_{i}})r}(1-e^{-\lambda r})e^{-r}=:a_{\lambda,r,v},$
where the first factor is the probability that the edge $\{\rho,x_{i}\}$ is open, the second is the probability that the edge is not updated until time $r$, the third is the probability that $\rho$ infects $x_{i}$ before time $r$, and the last one is the probability that the infection at $x_{i}$ remains until time $r$. We observe that we have the same lower bound for $\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(x_{i}\in\mathbf{C}_{r})$ since the kernel $p(\cdot,\cdot)$ and the speed function $v(\cdot,\cdot)$ are symmetric. Then, putting everything together in (5.3) and taking $\liminf_{t\to\infty}$ yields
$\liminf_{t\to\infty}\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(\rho\in\mathbf{C}_{t})$ $\geq a_{r,\lambda,v}^{2}\liminf_{t\to\infty}\mathbb{P}_{\mathcal{T}_{i}}^{\{x_{i}\}}(x_{i}\in\mathbf{C}_{t-2r}).$
In other words, we have deduced that
$\beta_{\mathcal{T}}(\lambda,v,\{\rho\})\geq a_{\lambda,r,v}^{2}\beta_{\mathcal{T}_{i}}(\lambda,v,\{x_{i}\}).$
Since for $\lambda,r,v>0$, we have $a_{\lambda,r,v}>0$, it follows that if $\beta_{\mathcal{T}_{i}}(\lambda,v,\{x_{i}\})>0$ for some $i\in\{1,\dots k\}$, then $\beta_{\mathcal{T}}(\lambda,v,\{\rho\})>0$. This proves that $\lambda_{2}(\mathcal{T})$ is constant on the event $\{|\mathcal{T}|=\infty\}$. For the last claim we remark that it holds that
$\mathbb{P}^{\{\rho\}}(\mathbf{C}_{t}\neq\emptyset\,\forall\,t\geq 0)$ $=\mathbb{E}\big{[}\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(\mathbf{C}_{t}\neq\emptyset\,\forall\,t\geq 0)\big{]}=\mathbb{E}\big{[}\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(\mathbf{C}_{t}\neq\emptyset\,\forall\,t\geq 0)\mathds{1}_{\{|\mathcal{T}|=\infty\}}\big{]}$
$\liminf_{t\to\infty}\mathbb{P}^{\{\rho\}}(\rho\in\mathbf{C}_{t})$ $=\mathbb{E}\big{[}\liminf_{t\to\infty}\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(\rho\in\mathbf{C}_{t})\big{]}=\mathbb{E}\big{[}\big{(}\liminf_{t\to\infty}\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(\rho\in\mathbf{C}_{t})\big{)}\mathds{1}_{\{|\mathcal{T}|=\infty\}}\big{]}.$
Since $\lambda_{1}(\mathcal{T})$ and $\lambda_{2}(\mathcal{T})$ are constant on $\{|\mathcal{T}|=\infty\}$ it follows that
@@
$\mathbb{P}^{\{\rho\}}(\mathbf{C}_{t}\neq\emptyset\,\forall\,t\geq 0)>0\quad$ $\Leftrightarrow\quad\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(\mathbf{C}_{t}\neq\emptyset\,\forall\,t\geq 0)>0\,\,\text{ on }\,\,\{|\mathcal{T}|=\infty\}$
@@
$\liminf_{t\to\infty}\mathbb{P}^{\{\rho\}}(\rho\in\mathbf{C}_{t})>0\quad$ $\Leftrightarrow\quad\liminf_{t\to\infty}\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(\rho\in\mathbf{C}_{t})>0\,\,\text{ on }\,\,\{|\mathcal{T}|=\infty\}$
Thus, we can conclude that on the event $\{|\mathcal{T}|=\infty\}$ it holds that $\lambda_{1}(\mathcal{T})=\lambda_{1}$ and $\lambda_{2}(\mathcal{T})=\lambda_{2}$. ∎
@@ 6 Proofs in the general graph setting
In this section we prove Theorem 3.1, Corollary 3.3, and Propositions 3.3 and 3.5. In order to show Theorem 3.1, we will adapt the technique used in the proof of [11, Theorem 5]. Note that [11] always work on finite graphs while we are dealing with an infinite graph. Thus, we need to adjust the proof by using an approximation argument via finite subgraphs. Let $\mathbf{X}=(\mathbf{C}^{\mathbf{X}},\mathbf{B}^{\mathbf{X}})=(\mathbf{C}_{t}^{\mathbf{X}},\mathbf{B}_{t}^{\mathbf{X}})_{t\geq 0}$ be a process with values in $\mathcal{P}(V)\times\mathcal{P}(E)$, where we again call a vertex infected or healthy if $x\in\mathbf{C}^{\mathbf{X}}$ or respectively $x\notin\mathbf{C}^{\mathbf{X}}$. Furthermore we call an edge $\{x,y\}$ revealed if $\{x,y\}\in\mathbf{B}^{\mathbf{X}}$ and otherwise unrevealed.
@@
We will define a process $(\mathbf{C}^{\mathbf{X}},\mathbf{B}^{\mathbf{X}})$ that is coupled to the processes $(\mathbf{C},\mathbf{B})$ via the graphical representation using the same Poisson point processes $\Delta^{\inf}$, $\Delta^{\text{rec}}$,$\Delta^{\text{op}}$ and $\Delta^{\text{cl}}$, and then show that the process has the correct dynamics. The evolution of $\mathbf{C}_{t}^{\mathbf{X}}$ works in the same way as that of $\mathbf{C}_{t}$, the only difference being that a potential infection event $t\in\Delta^{\inf}_{e}$ can be used for infection if $e\in\mathbf{B}_{t}^{\mathbf{X}}$ is a revealed edge even if $e\notin\mathbf{B}_{t-}$ (so the edge $e$ is not open). Recovery events in $\Delta^{\text{rec}}$ are used by $\mathbf{C}$ and $\mathbf{C}^{\mathbf{X}}$ in the same way. We will reveal the edge $e=\{x,y\}$ at any potential infection event $t\in\Delta^{\inf}_{e}$ for which $e\in\mathbf{B}_{t-}$ and either $x\in\mathbf{C}_{t-}^{\mathbf{X}}$ or $y\in\mathbf{C}_{t-}^{\mathbf{X}}$. In this case the infection is also passed on along this edge at time $t$ for the process $\mathbf{C}^{\mathbf{X}}$ (since then $e\in\mathbf{B}_{t}^{\mathbf{X}}$). An edge will be unrevealed whenever it is updated, so at times $t\in\Delta^{\text{op}}\cup\Delta^{\text{cl}}$.
@@
From this construction, it is clear that we have $\mathbf{C}_{t}\subset\mathbf{C}_{t}^{\mathbf{X}}$ since the initial states are identical and all infection paths that are used by the process $\mathbf{C}$ can also be used by $\mathbf{C}^{\mathbf{X}}$. For this note that at any time $t\in\Delta^{\inf}$ an edge that is unrevealed but open at $t-$ will be revealed at time $t$ whenever it is used by an infection path for $\mathbf{C}$ and can then, as a revealed edge, also be used by an infection path for $\mathbf{C}^{\mathbf{X}}$. Also, revealed edges (whether open or closed) can at any $t\in\Delta^{\inf}$ be used by an infection path for $\mathbf{C}^{\mathbf{X}}$. Lastly, both processes use the same recovery events.
@@
Thus, it just remains to verify that the process $(\mathbf{C}^{\mathbf{X}},\mathbf{B}^{\mathbf{X}})$ we defined has the dynamics that was described in (i) to (iv) before the statement of Lemma 6.1. The points (ii)-(iv) follow immediately: For (ii) we simply observe that the rate $\lambda$ is a consequence of the fact that $\Delta^{\inf}$ is for every edge constructed from a Poisson process at rate $\lambda$. Likewise, (iii) follows since update events in $\Delta^{\text{op}}\cup\Delta^{\text{cl}}$ for any edge coincide with unrevealment events and happen at rate $v(d_{x},d_{y})$. Finally, recoveries use $\Delta^{\text{rec}}$, which comes from a rate 1 Poisson process at every vertex, and this implies (iv).
Only the rate claimed in (i) takes a bit more deliberation. First, we note that an edge $e=\{x,y\}$ is only unrevealed at a time $t\in\Delta^{\inf}_{e}$ if the last event that affected $e$ prior to $t$ was an update event (as opposed to $e$ being revealed and used by an infection path). The outcome of this update event, an open edge with probability $p(d_{x},d_{y})$ and a closed edge otherwise, has thus not been relevant for both infection processes $\bf C$ and $\bf C^{\mathbf{X}}$ before time $t$. Thus, $\mathds{1}_{\{e\in\mathbf{B}_{t}\}}$ is a $\text{Ber}(p(d_{x},d_{y}))$ random variable whose value is independent of $\Delta^{\inf}_{e}$ and independent of other such Bernoulli random variables (defined analogously for unrevealed edges at infection times). Thus, the rate $\lambda p(d_{x},d_{y})$ in (i) for revealment (and potential infection) follows from thinning the rate $\lambda$ Poisson process $\Delta^{\inf}_{e}$ with independent Ber($p(d_{x},d_{y})$) random variables. ∎
@@
In order to be able to apply the generator to the function $f$, we define the following approximating process. For $n\in\mathbb{N}$, set $A_{n}:=(V_{n},E_{n})$ with
@@
$\mathcal{A}_{n}f(X)\leq$ $\sum_{x\in V_{n}}W(d_{x})\Big{(}\lambda K\Big{(}1+\frac{2\lambda}{\underline{v}^{2}}\Big{)}+\sum_{\{x,y\}\notin B^{X}}\frac{2\lambda^{2}}{v(d_{x},d_{y})^{2}}p(d_{x},d_{y})-\Big{(}\frac{\underline{v}}{2}\wedge 1\Big{)}\Big{)}H_{x}(X).$
@@
      An infection event between $x$ and one of its neighbors $y$ occurs and the edge $\{x,y\}$ was previously unrevealed, i.e. $\{x,y\}\notin B^{X}$. This would cause a change of $\tfrac{2\lambda}{v(d_{x},d_{y})^{2}}$. Now summing over the neighbours of $x$ yields that the total rate of change is
      $\sum_{\{x,y\}\notin B^{X}}\lambda p(d_{x},d_{y})\frac{2\lambda}{v(d_{x},d_{y})^{2}}.$
@@
      The cumulative change caused by infection events along unrevealed edges between $x$ and its neighbors is
@@
      If the infection event between $x$ and its neighbors happens at a revealed edge the cumulative change is
@@
      The cumulative change of all update events at revealed edges between $x$ and its neighbors is
@@
$\mathcal{A}_{n}f(X)=$ $\sum_{x\in C^{X}}W(d_{x})\bigg{(}\bigg{(}\sum_{\{x,y\}\notin B^{X}}\lambda p(d_{x},d_{y})\frac{2\lambda}{v(d_{x},d_{y})^{2}}\bigg{)}-2R_{x}(X)+\big{(}R_{x}(X)-1\big{)}\bigg{)}$
$+\sum_{x\notin C^{X}}W(d_{x})\bigg{(}\bigg{(}\sum_{y\in C^{X},\{x,y\}\in B^{X}}\lambda\big{(}1-R_{x}(X)\big{)}\bigg{)}-\lambda S_{x}(X)$
$-2R_{x}(X)+\sum_{y\in C^{X},\{x,y\}\notin B^{X}}\lambda p(d_{x},d_{y})\bigg{(}1-R_{x}(X)+\frac{2\lambda}{v(d_{x},d_{y})^{2}}\bigg{)}\bigg{)}.$
(Note that the sum over $x\notin C^{X}$ is only over $x\in V_{n}$ which are not infected. We can simplify and ignore the negative contributions from $\big{(}\sum_{y\in C^{X},\{x,y\}\in B^{X}}\lambda\big{(}1-R_{x}(X)\big{)}\big{)}-\lambda S_{x}(X)$ in the second and $-R_{x}(X)$ in the very last sum in order to obtain
$\mathcal{A}_{n}f(X)\leq$ $\sum_{x\in C^{X}}W(d_{x})\bigg{(}\bigg{(}\sum_{\{x,y\}\notin B^{X}}\frac{2\lambda^{2}}{v(d_{x},d_{y})^{2}}p(d_{x},d_{y})\bigg{)}-1-R_{x}(X)\bigg{)}$ (6.3)
@@
$\mathcal{A}_{n}f(X)\leq$ $\sum_{x\in C^{X}}\lambda K\Big{(}1+\frac{2\lambda}{\underline{v}^{2}}\Big{)}W(d_{x})+\sum_{x\in C^{X}}W(d_{x})\sum_{\{x,y\}\notin B^{X}}\frac{2\lambda^{2}}{v(d_{x},d_{y})^{2}}p(d_{x},d_{y})$
@@
$\mathcal{A}_{n}f(X)\leq$ $\sum_{x\in V_{n}}W(d_{x})\Big{(}\lambda K\Big{(}1+\frac{2\lambda}{\underline{v}^{2}}\Big{)}+\sum_{\{x,y\}\notin B^{X}}\frac{2\lambda^{2}}{v(d_{x},d_{y})^{2}}p(d_{x},d_{y})-\Big{(}\frac{\underline{v}}{2}\wedge 1\Big{)}\Big{)}H_{x}(X),$
@@
$\vartheta:=\lambda K\Big{(}1+\frac{2\lambda}{\underline{v}^{2}}\Big{)}+\frac{2\lambda^{2}}{\underline{v}^{2}}\widetilde{K}-\Big{(}\frac{\underline{v}}{2}\wedge 1\Big{)}<0.$ (6.6)
Then it follows that the process $M^{n}=(M_{t}^{n})_{t\geq 0}$ defined as in (6.1) is a supermartingale for every $n\in\mathbb{N}$, which converges almost surely to 00 as $t\to\infty$. This implies in particular that $\mathbf{C}$ goes extinct almost surely.
@@
$\partial_{t}T^{n}_{t}f(X)=T^{n}_{t}\mathcal{A}_{n}f(X)\leq T^{n}_{t}\Big{(}\lambda K\Big{(}1+\frac{2\lambda}{\underline{v}^{2}}\Big{)}+\frac{2\lambda^{2}}{\underline{v}^{2}}\widetilde{K}-\Big{(}\frac{\underline{v}}{2}\wedge 1\Big{)}\Big{)}f(X)=\vartheta T^{n}_{t}f(X).$

### v2

@@ 5.2 Properties of $(\mathbf{C},\mathbf{B})$ on $\mathcal{T}$
Let us now prove the statement for $\lambda_{2}(\mathcal{T})$. Denote by
$\beta_{\mathcal{T}}(\lambda,v,\{\rho\}):=\mathbb{P}_{\mathcal{T}}^{\{\rho\}}(\forall s\geq 0\,\exists t\geq s\,:\,\rho\in\mathbf{C}_{t}).$
The same reasoning as in the argument for $\lambda_{1}(\mathcal{T})$ applies, with the only difference that we need to argue that conditionally on $\mathcal{T}$, if $\beta_{\mathcal{T}_{i}}(\lambda,v,\{x_{i}\})>0$ for some $i\in\{1,\dots,k\}$, then $\beta_{\mathcal{T}}(\lambda,v,\{\rho\})>0$. First of all, $\rho$ infects $x_{i}$ in finite time with positive probability. Then, by assumption (with positive probability) $x_{i}$ is infected at times $t_{j}\rightarrow\infty$ as $j\to\infty$ in the process restricted to $\mathcal{T}_{i}$. By moving to a subsequence (and the Borel-Cantelli lemma), we may assume that $|t_{j+1}-t_{j}|\geq 1$ and also that $x_{i}$ remains infected during $t_{j},t_{j}+1]$. Then, let $E_{j}$ be the event that there is an infection event in $t_{j},t_{j}+1]$ along the edge $\{x_{i},\rho\}$ in the graphical construction (restricted to this edge) when the edge is open. Conditionally on the process restricted to $\mathcal{T}_{i}$, the events $E_{j}$ are independent and have positive probability. Thus, the conditional probability that infinitely many of the events $E_{j}$ happen is equal to $1$ by the Borel-Cantelli lemma. If this happens then $\rho\in\mathbf{C}_{s_{i}}$ for some $s_{i}\rightarrow\infty$. Therefore, $\beta_{\mathcal{T}}(\lambda,v,\{\rho\})>0$.
Note that as the CPDG dies out almost surely on a finite tree, we have that
$\mathbb{P}^{\{\rho\}}(\mathbf{C}_{t}\neq\emptyset\,\forall\,t\geq 0)$ $=\mathbb{E}\big[\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(\mathbf{C}_{t}\neq\emptyset\,\forall\,t\geq 0)\big]=\mathbb{E}\big[\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(\mathbf{C}_{t}\neq\emptyset\,\forall\,t\geq 0)\mathds{1}_{\{|\mathcal{T}|=\infty\}}\big],$
and similarly
$\mathbb{P}^{\{\rho\}}(\forall s\geq 0\,:\,\exists t\geq s\,:\,\rho\in\mathbf{C}_{t})$ $=\mathbb{E}\big[\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(\forall s\geq 0\,:\,\exists t\geq s\,:\,\rho\in\mathbf{C}_{t})\mathds{1}_{\{|\mathcal{T}|=\infty\}}\big].$
Thus, by (5.3) (and the analogous expression for $\beta_{\mathcal{T}}$)
@@
$\mathbb{P}^{\{\rho\}}(\mathbf{C}_{t}\neq\emptyset\,\forall\,t\geq 0)>0\quad$ $\Leftrightarrow\quad\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(\mathbf{C}_{t}\neq\emptyset\,\forall\,t\geq 0)>0\,\,\text{almost surely on }\,\,\{|\mathcal{T}|=\infty\},$
@@
$\mathbb{P}^{\{\rho\}}(\forall s\geq 0\,:\,\exists t\geq s\,:\,\rho\in\mathbf{C}_{t})>0\quad$ $\Leftrightarrow\quad\mathbb{P}^{\{\rho\}}_{\mathcal{T}}(\forall s\geq 0\,:\,\exists t\geq s\,:\,\rho\in\mathbf{C}_{t})>0\,\,\text{almost surely on }\{|\mathcal{T}|=\infty\}.$
This proves the statement of the proposition. ∎
@@ 6 Proofs in the general graph setting
In this section we prove Theorem 3.1, Corollary 3.3, and Propositions 3.3 and 3.5. In order to show Theorem 3.1, we will adapt the technique used in the proof of [11, Theorem 5]. Note that [11] always work on finite graphs while we are dealing with an infinite graph. Thus, we need to adjust the proof by using an approximation argument via finite subgraphs. Let $\mathbf{X}=(\mathbf{C}^{\mathbf{X}},\mathbf{B}^{\mathbf{X}})=(\mathbf{C}_{t}^{\mathbf{X}},\mathbf{B}_{t}^{\mathbf{X}})_{t\geq 0}$ be a process with values in $\mathcal{P}(V)\times\mathcal{P}(E)$, where we again call a vertex infected or healthy if $x\in\mathbf{C}^{\mathbf{X}}$ or respectively $x\notin\mathbf{C}^{\mathbf{X}}$. Furthermore we call an edge $\{x,y\}$ revealed if $\{x,y\}\in\mathbf{B}^{\mathbf{X}}$ and otherwise unrevealed. This auxiliary process obeys the following rules:
@@
We will define a process $(\mathbf{C}^{\mathbf{X}},\mathbf{B}^{\mathbf{X}})$ that is coupled to the processes $(\mathbf{C},\mathbf{B})$ via the graphical representation using the same Poisson point processes $\Delta^{\inf}$, $\Delta^{\text{rec}}$, $\Delta^{\text{op}}$, $\Delta^{\text{cl}}$, as well as some additional independent randomness, and then show that the process has the correct dynamics. In order to prepare for the coupling we first define a thinned version of $\Delta^{\inf}$ that will provide us with the needed additional independent randomness. For each $e=\{x,y\}$ we use an independent sequence of i.i.d. $\text{Ber}(p(d_{x},d_{y}))$ random variables to thin $\Delta^{\inf}_{e}$ so that we obtain a Poisson point process $\Delta^{\inf,p}_{e}$ with rate $\lambda p(d_{x},d_{y}).$
The evolution of $\mathbf{C}_{t}^{\mathbf{X}}$ is defined as follows: If $e\in\mathbf{B}_{t-}^{\mathbf{X}}$ such that $e$ is a revealed edge right before time $t$ then a potential infection event $t\in\Delta^{\inf}_{e}$ can be used for infection. If $e\notin\mathbf{B}_{t-}^{\mathbf{X}}$ such that $e$ is an unrevealed edge right before time $t$ then at a potential infection event $t\in\Delta^{\inf}_{e}$ the edge $e=\{x,y\}$ can be revealed (and infection can occur along the edge) if $x\in\mathbf{C}_{t-}^{\mathbf{X}}$ or $y\in\mathbf{C}_{t-}^{\mathbf{X}}$. Whether this revealment and infection actually occurs (otherwise nothing happens) is decided by independent $\text{Ber}(p(d_{x},d_{y}))$ random variables. As a consequence the overall rate of revealment and infection in this setting is then $\lambda p(d_{x},d_{y}).$ Generally, in order to assure that $\mathbf{C}_{t}\subset\mathbf{C}_{t}^{\mathbf{X}}$ for all $t\geq 0$ we would like to use $\mathds{1}_{\{e\in\mathbf{B}_{t-}\}}$, thus revealment and infection happens if the edge $e$ is open. However, if the edge is closed, and there is no revealment then at the next potential revealment (if it occurs before the next update event) we cannot use the same Bernoulli random variable but need another independent Bernoulli random variable. In this case we use the independent Bernoulli random variables that defined the thinning of $\Delta^{\inf}_{e}$ (formally we use $\mathds{1}_{\{t\in\Delta^{\inf,p}_{e}\}}$). This then assures independence of the $\text{Ber}(p(d_{x},d_{y}))$ random variables. Notice that using this independent randomness to decide whether revealment and infection happens does not disturb the monotone coupling as we can in this situation only have additional infections for $\mathbf{C}_{t}^{\mathbf{X}}$ since the edge is closed and no infections occur for $\mathbf{C}_{t}$ until the next update of the edge. (It is in fact only due to such revealment and infection events that we may obtain more infected sites in $\mathbf{C}^{\mathbf{X}}$ than in $\mathbf{C}$.) Furthermore, an edge will be unrevealed whenever it is updated, so at times $t\in\Delta^{\text{op}}\cup\Delta^{\text{cl}}$. Also, the recovery events in $\Delta^{\text{rec}}$ are used by $\mathbf{C}$ and $\mathbf{C}^{\mathbf{X}}$ in the same way.
@@
From this construction, it is clear that we have $\mathbf{C}_{t}\subset\mathbf{C}_{t}^{\mathbf{X}}$ since the initial states are identical and all infection paths that are used by the process $\mathbf{C}$ can also be used by $\mathbf{C}^{\mathbf{X}}$. For this note that at any time $t\in\Delta^{\inf}$ an edge that is unrevealed but open at $t$- will be revealed at time $t$ whenever it is used by an infection path for $\mathbf{C}$ and that we then also have an infection path for $\mathbf{C}^{\mathbf{X}}$. Also, revealed edges (whether open or closed) can at any $t\in\Delta^{\inf}$ be used by an infection path for $\mathbf{C}^{\mathbf{X}}$. Lastly, both processes use the same recovery events.
@@
Thus, it just remains to verify that the process $(\mathbf{C}^{\mathbf{X}},\mathbf{B}^{\mathbf{X}})$ we defined has the dynamics that was described in (i) to (iv) before the statement of Lemma 6.1. We have already argued in the construction above that revealment happens for edges $e=\{x,y\}$ at rate $\lambda p(d_{x},d_{y})$ while $x\in\mathbf{C}^{\mathbf{X}}$ or $y\in\mathbf{C}^{\mathbf{X}}$. The points (ii)-(iv) follow immediately: For (ii) we simply observe that the rate $\lambda$ is a consequence of the fact that $\Delta^{\inf}$ is for every edge constructed from a Poisson process at rate $\lambda$. Likewise, (iii) follows since update events in $\Delta^{\text{op}}\cup\Delta^{\text{cl}}$ for any edge coincide with unrevealment events and happen at rate $v(d_{x},d_{y})$. Finally, recoveries use $\Delta^{\text{rec}}$, which comes from a rate 1 Poisson process at every vertex, and this implies (iv). ∎
@@
In order to be able to apply the generator to the function $f$, we define the following approximating process. For $n\in\mathbb{N}$, set $A_{n}:=V_{n}\times E_{n}$ with
@@
$\mathcal{A}_{n}f(X)\leq$ $\sum_{x\in V_{n}}W(d_{x})\Big(\lambda K\Big(1+\frac{2\lambda}{\underline{v}^{2}}\Big)+\sum_{\{x,y\}\notin B^{X}}\frac{4\lambda^{2}}{v(d_{x},d_{y})^{2}}p(d_{x},d_{y})-\Big(\frac{\underline{v}}{2}\wedge 1\Big)\Big)H_{x}(X).$
@@
      Either $x$ attempts to infect one of its neighbours $y$ and $\{x,y\}$ changes from being unrevealed to revealed or one of the infected neighbours $y$ of $x$ attempts to infect $x$ along a previously unrevealed edge that now becomes revealed. Each of these events causes a change of $\tfrac{2\lambda}{v(d_{x},d_{y})^{2}}$, so that the cumulative change is
      $\sum_{y\in V_{n}\,:\,\{x,y\}\notin B^{X}}\lambda p(d_{x},d_{y})\frac{2\lambda}{v(d_{x},d_{y})^{2}}+\sum_{y\in C^{X}\,:\,\{x,y\}\notin B^{X}}\lambda p(d_{x},d_{y})\frac{2\lambda}{v(d_{x},d_{y})^{2}}.$
@@
      The cumulative change caused by infection events along unrevealed edges between $x$ and its neighbours is
@@
      If the infection event between $x$ and its neighbours happens at a revealed edge the cumulative change is
@@
      The cumulative change of all update events at revealed edges between $x$ and its neighbours is
@@
$\mathcal{A}_{n}f(X)=$ $\sum_{x\in C^{X}}W(d_{x})\bigg(\sum_{\{x,y\}\notin B^{X}}\frac{2\lambda^{2}p(d_{x},d_{y})}{v(d_{x},d_{y})^{2}}+\sum_{y\in C^{X}:\{x,y\}\notin B^{X}}\frac{2\lambda^{2}p(d_{x},d_{y})}{v(d_{x},d_{y})^{2}}-2R_{x}(X)+\big(R_{x}(X)-1\big)\bigg)$
$+\sum_{x\notin C^{X}}W(d_{x})\bigg(\sum_{y\in C^{X},\{x,y\}\in B^{X}}\lambda\big(1-R_{x}(X)\big)-\lambda S_{x}(X)$
$\phantom{XXXXXXXX}-2R_{x}(X)+\sum_{y\in C^{X},\{x,y\}\notin B^{X}}\lambda p(d_{x},d_{y})\bigg(1-R_{x}(X)+\frac{2\lambda}{v(d_{x},d_{y})^{2}}\bigg)\bigg).$
Note that the sum over $x\notin C^{X}$ is only over $x\in V_{n}$ which are not infected. We can simplify and ignore the negative contributions from $\big(\sum_{y\in C^{X},\{x,y\}\in B^{X}}\lambda\big(1-R_{x}(X)\big)\big)-\lambda S_{x}(X)$ in the second and $R_{x}(X)$ in the very last sum in order to obtain
$\mathcal{A}_{n}f(X)\leq$ $\sum_{x\in C^{X}}W(d_{x})\bigg(\sum_{\{x,y\}\notin B^{X}}\frac{4\lambda^{2}}{v(d_{x},d_{y})^{2}}p(d_{x},d_{y})-1-R_{x}(X)\bigg)$ (6.3)
@@
$\mathcal{A}_{n}f(X)\leq$ $\sum_{x\in C^{X}}\lambda K\Big(1+\frac{2\lambda}{\underline{v}^{2}}\Big)W(d_{x})+\sum_{x\in C^{X}}W(d_{x})\sum_{\{x,y\}\notin B^{X}}\frac{4\lambda^{2}}{v(d_{x},d_{y})^{2}}p(d_{x},d_{y})$
@@
$\mathcal{A}_{n}f(X)\leq$ $\sum_{x\in V_{n}}W(d_{x})\Big(\lambda K\Big(1+\frac{2\lambda}{\underline{v}^{2}}\Big)+\sum_{\{x,y\}\notin B^{X}}\frac{4\lambda^{2}}{v(d_{x},d_{y})^{2}}p(d_{x},d_{y})-\Big(\frac{\underline{v}}{2}\wedge 1\Big)\Big)H_{x}(X),$
@@
$\vartheta:=\lambda K\Big(1+\frac{2\lambda}{\underline{v}^{2}}\Big)+4\lambda^{2}K-\Big(\frac{\underline{v}}{2}\wedge 1\Big)<0.$ (6.6)
Then it follows that the process $M^{n}=(M_{t}^{n})_{t\geq 0}$ defined as in (6.1) is a supermartingale for every $n\in\mathbb{N}$, which converges almost surely to 0 as $t\to\infty$. This implies in particular that $\mathbf{C}$ goes extinct almost surely.
@@
$\partial_{t}T^{n}_{t}f(X)=T^{n}_{t}\mathcal{A}_{n}f(X)\leq T^{n}_{t}\Big(\lambda K\Big(1+\frac{2\lambda}{\underline{v}^{2}}\Big)+4\lambda^{2}K-\Big(\frac{\underline{v}}{2}\wedge 1\Big)\Big)f(X)=\vartheta T^{n}_{t}f(X).$